OpenStudy (studygurl14):

NEXT ONE> PAY ATTENTION> RULES ARE DIFFERENT! MDEAL TO FIRST PERSON WHO SHOWS ALL THE STEPS AND ANSWERS CORRECTLY!

OpenStudy (anonymous):

lol

OpenStudy (studygurl14):

Solve for x: \[x^{3} - 2 = 25\]

OpenStudy (studygurl14):

Any takers?

OpenStudy (anonymous):

3^3-2=25

OpenStudy (studygurl14):

steps please @DragonFlame...tho the answer is correct

OpenStudy (anonymous):

Whoops nvm I messed up

OpenStudy (studygurl14):

Okay fine @DragonFlame I'll give it to you...you did good

OpenStudy (anonymous):

Should I still put in how I did it?

OpenStudy (studygurl14):

if you want

OpenStudy (anonymous):

x^3 - 2 = 25 add 2 to both sides x^3 = 25 + 2 x^3 = 27 x = cubed root of 27 x = 3